OpenStudy (anonymous):

Determine whether the Nonhomogenous system is consistent. 1 3 10 18 -2 7 32 29 -1 3 14 12 1 1 2 8

OpenStudy (anonymous):

I don't want sit here for free. give you the answer: the system is consistent under the condition of \[\huge\x_1=3+2x_3\] \[\huge\color{red}{x_2=5-4x_3}\] \[\huge\color{purple}{x_3~free}\]